Determine the quadratic equation with integral coefficients having the roots:
3/2 and -1/2

Answer by shweta(56)

Equation of quadratic equation:
x^2 -(sum of the roots)x + product of the roots
sum of the roots: 3/2-1/2=2/2=1
product of the roots= (-1/2)*(3/2)= -3/4
equation of the quadratic equation:
x^2-x-3/4=0
4x^2-4x-3=0